Q1- Deni shuffles the playing cards and a heart comes out of the deck 1 out of 13 draws and X is the random variable "the sum of the numbers that turn up". Find the probability distribution of X.

Q2- Charlie shuffles the playing cards and a queen come out of the deck only once out of 4 times. X is the random variable "the sum of the numbers that turn up". Find the probability distribution of X.

Q3-In a box of 15 chocolates. 3 are dark chocolate the rest are milk chocolate. You will be randomly selecting 4 chocolates from the box. Let X be the number of dark chocolates. Find the probability distribution of X.
Answer #1

Since Q1 and Q2 are not clear, I am helping you with Q3 ...

n = 15, p = 3/15 = 0.2, q = 1 - p = 0.8, x = 4
Probability distribution of the number of dark chocolates:
x P(x) = C(n, x) p^x q^(n - x)
0 0.0352
1 0.1319
2 0.2309
3 0.2501
4 0.1876